OpenStudy (anonymous):

The area of a circle with a radius of 15 in., to the nearest tenth is _____. Be sure to include units

OpenStudy (paxpolaris):

area of a circle \[\large=\pi r^2\]

OpenStudy (anonymous):

225*3.14

OpenStudy (paxpolaris):

\[=3.14\times (15in.)^2\\=3.14\times225\ in^2\]

OpenStudy (anonymous):

thanks all